OpenStudy (ericamarie7):

Someone help me with this Binomial please. (-2z^2-14)-(-12z^2)

OpenStudy (badsheep):

Simplifying (-2z2 + -14) + -1(-12z2) = 0 Reorder the terms: (-14 + -2z2) + -1(-12z2) = 0 Remove parenthesis around (-14 + -2z2) -14 + -2z2 + -1(-12z2) = 0 Remove parenthesis around (-12z2) -14 + -2z2 + -1 * -12z2 = 0 Multiply -1 * -12 -14 + -2z2 + 12z2 = 0 Combine like terms: -2z2 + 12z2 = 10z2 -14 + 10z2 = 0 Solving -14 + 10z2 = 0 Solving for variable 'z'. Move all terms containing z to the left, all other terms to the right. Add '14' to each side of the equation. -14 + 14 + 10z2 = 0 + 14 Combine like terms: -14 + 14 = 0 0 + 10z2 = 0 + 14 10z2 = 0 + 14 Combine like terms: 0 + 14 = 14 10z2 = 14 Divide each side by '10'. z2 = 1.4 Simplifying z2 = 1.4 Take the square root of each side: z = {-1.183215957, 1.183215957}